# Seek-CAD: A Self-refined Generative Modeling for 3D Parametric CAD Using Local Inference via DeepSeek

### v2

@@ A.2 CAD dataset Based on SSR Paradigm
Figure 7: Qualitative comparison between our dataset and the DeepCAD [54] dataset. Compared to DeepCAD, our dataset captures more realistic and structurally complex industrial designs, supporting a broader range of modeling operations such as spline, revolve, chamfer, fillet, and shell. Our extended dataset also features richer geometric details and greater diversity in modeling strategies.
Figure 8: Statistical comparison between our complete dataset and the DeepCAD [54] dataset, which also is a subset of our dataset. The left plot illustrates the distribution of feature lengths, while the right plot shows the distribution of average curve lengths.
@@ A.3 CAD Code Representation
{python}
class Loop: def __init__(self) -> None: … # Define start point of this loop def moveTo(self, x: float, y: float) -> Self: … # Draw straight line to point def lineTo(self, x: float, y: float) -> Self: … # Arc to p2 via p1 def threePointArc(self, p1: Tuple[float, float], p2: Tuple[float, float]) -> Self: … # Spline through given points def splineTo(self, *p: Tuple[float, float]) -> Self: … # Close loop (make a line back to start point) def close(self) -> Self: … # Draw circle with center at current point def circle(self, radius: float) -> Self: … # assign tag to the current point def pointTag(self, tag: str) -> Self: … # assign tag to the current curve def curveTag(self, tag: str) -> Self: …
class Profile: # Create a face with a optional tag for reference def __init__(self, tag: Optional[str] = None) -> None: … # Add one or more loops to the face def addLoop(self, *loops: Loop) -> None: …
class Sketch: # plane format: "origin": [x, y, z], "x_axis": [x, y, z], "normal": [x, y, z] def __init__(self, plane: Dict) -> None: … def addProfile(self, *profiles: Profile) -> None: …
class CADShape(ABC): def __init__(self) -> None: … # entities: list of dicts, each with "capType": "START"|"END"|"SWEEP", ’referenceId’: str, specifying referenced entities def Chamfer(self, width: float, entities: List[Dict]) -> Self: … def Fillet(self, radius: float, entities: List[Dict]) -> Self: … def Shell(self, thickness: float, entities: List[Dict]) -> Self: … def union(self, shape: CADShape) -> Self: … def cut(self, shape: CADShape) -> Self: … def intersect(self, shape: CADShape) -> Self: …
class Extrude(CADShape): def __init__(self, sketch: Sketch, distance: Union[float, Tuple[float, float]]) -> None: …
class Revolve(CADShape): # axis format: "point": [x, y, z], "direction": [x, y, z], axis is defined by a point and a direction def __init__(self, sketch: Sketch, axis: Dict, angle: Union[float, Tuple[float, float]]) -> None: …
@@ A.4 RAG Settings
To equip the SSR-based CAD corpus locally for the retrieval-augmented generation of Seek-CAD, we use Dify [6] in 0.15.3 version, which is a docker-based [7] platform. For each query, we use bge-m3 [5] to embed it into a vector that is used to calculate similarity score with CAD models stored in the local SSR-based corpus. Practically, we select Top-3 samples based on a hybrid searching strategy (Sec 3.1) for the augmentation retrieval. To incorporate the retrieved content into the input for Seek-CAD, we apply a chunk-based prompt construction strategy. Each retrieved pair $(d_{j},s_{j})\in\mathcal{R}_{T}$ is formatted as:
[…]
@@
$T\oplus(d_{j},c_{j})=\ \texttt{"Query: "}T\oplus\ \texttt{Chunk}_{1}\ \oplus\ \dots\ \oplus\ \texttt{Chunk}_{k}.$ (12)

### v3

@@ A.2 CAD dataset Based on SSR Paradigm
Figure 7: Qualitative comparison between our dataset and the DeepCAD Wu et al. (2021) dataset. Compared to DeepCAD, our dataset captures more realistic and structurally complex industrial designs, supporting a broader range of modeling operations such as spline, revolve, chamfer, fillet, and shell. Our extended dataset also features richer geometric details and greater diversity in modeling strategies.
Figure 8: Statistical comparison between our complete dataset and the DeepCAD Wu et al. (2021) dataset, which also is a subset of our dataset. The left plot illustrates the distribution of feature lengths, while the right plot shows the distribution of average curve lengths.
@@ A.3 CAD Code Representation
def __init__(self) -> None: ...
# Define start point of this loop
def moveTo(self, x: float, y: float) -> Self: ...
# Draw straight line to point
def lineTo(self, x: float, y: float) -> Self: ...
# Arc to p2 via p1
def threePointArc(self, p1: Tuple[float, float], p2: Tuple[float, float]) -> Self: ...
# Spline through given points
def splineTo(self, *p: Tuple[float, float]) -> Self: ...
# Close loop (make a line back to start point)
def close(self) -> Self: ...
# Draw circle with center at current point
def circle(self, radius: float) -> Self: ...
# assign tag to the current point
def pointTag(self, tag: str) -> Self: ...
# assign tag to the current curve
def curveTag(self, tag: str) -> Self: ...
class Profile:
# Create a face with a optional tag for reference
def __init__(self, tag: Optional[str] = None) -> None: ...
# Add one or more loops to the face
def addLoop(self, *loops: Loop) -> None: ...
class Sketch:
# plane format: {"origin": [x, y, z], "x_axis": [x, y, z], "normal": [x, y, z]}
def __init__(self, plane: Dict) -> None: ...
def addProfile(self, *profiles: Profile) -> None: ...
class CADShape(ABC):
def __init__(self) -> None: ...
# entities: list of dicts, each with {"capType": "START"|"END"|"SWEEP", ’referenceId’: str}, specifying referenced entities
def Chamfer(self, width: float, entities: List[Dict]) -> Self: ...
def Fillet(self, radius: float, entities: List[Dict]) -> Self: ...
def Shell(self, thickness: float, entities: List[Dict]) -> Self: ...
def union(self, shape: CADShape) -> Self: ...
def cut(self, shape: CADShape) -> Self: ...
def intersect(self, shape: CADShape) -> Self: ...
class Extrude(CADShape):
def __init__(self, sketch: Sketch, distance: Union[float, Tuple[float, float]]) -> None: ...
class Revolve(CADShape):
# axis format: {"point": [x, y, z], "direction": [x, y, z]}, axis is defined by a point and a direction
def __init__(self, sketch: Sketch, axis: Dict, angle: Union[float, Tuple[float, float]]) -> None: ...
@@ A.4 RAG Settings
To equip the SSR-based CAD corpus locally for the retrieval-augmented generation of Seek-CAD, we use Dify Dify in 0.15.3 version, which is a docker-based Docker platform. For each query, we use bge-m3 Chen et al. (2024) to embed it into a vector that is used to calculate similarity score with CAD models stored in the local SSR-based corpus. Practically, we select Top-3 samples based on a hybrid searching strategy (Sec 3.1) for the augmentation retrieval. To incorporate the retrieved content into the input for Seek-CAD, we apply a chunk-based prompt construction strategy. Each retrieved pair $(d_{j},s_{j})\in\mathcal{R}_{T}$ is formatted as:
$\texttt{[Chunk}_{j}\texttt{]}:=\texttt{"Description: "}d_{j}\ \|\ \texttt{"CAD code: "}s_{j},$ (11)
@@
$T\oplus(d_{j},s_{j})=\ \texttt{"Query: "}T\oplus\ \texttt{Chunk}_{1}\ \oplus\ \dots\ \oplus\ \texttt{Chunk}_{k}.\ \$(12)
@@ A.10 More showcases by Seek-CAD
A.11 The enlarged version of key design modules within Seek-CAD framework
Figure 15: The enlarged version of key design modules: (a) Initial CAD Code Generation. (b) CAD Code Refinement.